# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ D.2 Lower Bound for Stagger
Throughout the proof, denote by $\mathbf{E}_{\text{annotated}}$ the final set of states in $\mathbf{E}$ on which Stagger have requested expert annotations.
@@
By Lemma 21, if Stagger collects at most $\frac{HN_{0}}{12}$ interactive state-wise annotations as above, then with probability at least $\frac{1}{2}$, $|\mathbf{E}_{\text{annotated}}|$, the number of distinct states in $\mathbf{E}$ annotated is fewer than $N_{0}/3$. Consider a random rollout of $\hat{\pi}$, define the following events:
$F_{1}$ $:=\left\{s_{1}\notin\mathbf{E}_{\text{annotated}}\right\},$
@@
$\geq$ $\mathbb{E}_{\hat{\pi}}\left[H-\sum_{h=1}^{H}r_{h}\mid F_{1},F_{2}\right]P_{\hat{\pi}}(F_{1},F_{2})$
$\geq$ $\frac{2}{3}\cdot\frac{H-8}{H}\cdot\frac{10H-1}{10H}\cdot(H-1)\geq\frac{H}{2}$
Where we use our setting of $\beta=\frac{8}{H-8}$ and apply $H\geq 50$.
@@
$\Pr\left(|\mathbf{E}_{\text{annotated}}|\geq\frac{N_{0}}{3}\right)\leq\frac{1}{2}.$
@@
In this proof, we say that state $s$ is annotated at iteration $i$, if it has been annotated by the expert before iteration $i$ (excluding iteration $i$). Since each iteration of Stagger samples only one state uniformly from the current episode for annotation, we denote the indicator of expert annotating an unannotated state from $\mathbf{E}$ at iteration $i$ as $X_{i}\in\{0,1\}$. With this notation, we have the total number of annotated states by the end of iteration $N_{\mathrm{int}}$ as
$|\mathbf{E}_{\text{annotated}}|=\sum_{i=1}^{N_{\mathrm{int}}}X_{i}.$
Let $\mathcal{F}_{j}$ be the sigma-algebra generated by all information seen by Stagger up to iteration $j$. We now upper bound the expected value of $X_{i}$ conditioned on $\mathcal{F}_{i-1}$, for each $i$.
@@
$\mathbb{E}[|\mathbf{E}_{\text{annotated}}|]=\sum_{i=1}^{N_{\mathrm{int}}}\mathbb{E}[X_{i}]\leq\frac{2N_{\mathrm{int}}}{H}.$
@@
$\Pr(|\mathbf{E}_{\text{annotated}}|\geq N_{0}/3)\leq\frac{\mathbb{E}[X]}{N_{0}/3}\leq\frac{6N_{\mathrm{int}}}{HN_{0}}\leq\frac{1}{2}.$
@@ D.3 Upper Bound for Warm-Stagger
We divide the proof into four parts. In this proof, we denote by $\mathbf{E}_{\text{annotated}}$ and $\mathbf{E}_{\text{annotated}}^{\prime}$ the subsets of $\mathbf{E}$ and $\mathbf{E^{\prime}}$, respectively, that are annotated by the $N_{\mathrm{off}}$ offline expert demonstration trajectories.
@@
$F_{3}:=\{\mathbf{E}_{\text{annotated}}=\mathbf{E}\}.$
@@
, we define the event under which only a small fraction of $\mathbf{E^{\prime}}$ is covered by the expert. With $H\geq 50$, $H\geq\frac{5}{4}\log(10N_{0})$, $\beta=\frac{8}{H-8}$, $N_{1}\geq 500$, and $N_{1}\geq 160N_{0}$, the number of offline trajectories satisfies
@@
where we apply $H\geq\frac{5}{4}\log(10N_{0})$ and $\frac{H-16}{H-8}<\frac{4}{5}$ when $H\geq 50$. Thus, by Lemma 17,
$\mathbb{E}\!\left[|\mathbf{E}_{\text{annotated}}^{\prime}|\right]\leq\frac{N_{1}}{20}.$
@@
$\Pr\Big(|\mathbf{E}_{\text{annotated}}^{\prime}|>\frac{N_{1}}{4}-2\Big)\leq\Pr\Big(|\mathbf{E}_{\text{annotated}}^{\prime}|>\frac{N_{1}}{4}-\frac{2N_{1}}{500}\Big)\leq\frac{25}{121}<0.21.$
@@
$F_{4}:=\left\{|\mathbf{E}_{\text{annotated}}^{\prime}|\leq\frac{N_{1}}{4}-2\right\}.$
@@
, we show that under events $F_{3}$ and $F_{4}$, the interactive phase of Warm-Stagger annotates $\mathbf{b^{\prime}}$ with probability greater than 0.4 in each of the first three rollouts, such that $\mathbf{b^{\prime}}$ is annotated within three expert annotations with probability greater than 0.78.
Conditioned on $F_{3}$, the learner acts optimally on all states in $\mathbf{E}$. Conditioned on $F_{4}$, at least $\frac{3}{4}$ of the states in $\mathbf{E^{\prime}}$ are unannotated for each of the first three rollouts. Let $\tau$ be the first index with $s_{\tau}\in\mathbf{E^{\prime}}$. By Lemma 18,
$\Pr(\tau\leq H/5)\geq 0.79.$
Since $s_{\tau}$ is drawn uniformly from $\mathbf{E^{\prime}}$ and at most $N_{1}/4$ states are annotated under $B$,
$\Pr(s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime}\mid F_{3},F_{4},\tau\leq H/5)\;\geq\;\frac{3}{4}.$
@@
$\Pr\!\left(\pi(s_{\tau})\neq\pi^{E}(s_{\tau})\mid s_{\tau}\notin\mathbf{E}_{\text{annotated}}^{\prime},F_{3},F_{4},\tau\leq H/5\right)\geq 0.9,$
so with probability at least $0.9$ the learner transitions to $\mathbf{b^{\prime}}$ at time $\tau+1$ and stays there till the end of episode, accumulating no less than $0.8H$ $\mathbf{b^{\prime}}$ states with $H\geq 50$. Since Warm-Stagger samples one state uniformly from each rollout for annotation, this gives 0.8 probability of sampling $\mathbf{b^{\prime}}$ in this specific case.
Combining these factors, we obtain
[…]
@@
$F_{5}:=\{\mathbf{b^{\prime}}\text{ is annotated within 3 rollouts}\}.$
Conditioned on $F_{3}$ and $F_{4}$, the three rollouts has in each rollout the probability of sampling $\mathbf{b^{\prime}}$ is at least $0.4$. Hence
$\Pr(F_{5}\mid F_{3},F_{4})\;\geq\;1-(1-0.4)^{3}=1-0.6^{3}>0.78.$
@@
, on $F_{3}\cap F_{4}\cap F_{5}$, all states in $\mathbf{E}$ and the reset state $\mathbf{b^{\prime}}$ are annotated, under reward function $R_{1}$, the learner receives reward 1 at any step if it is in $\mathbf{E}$ or $\mathbf{E^{\prime}}$ or it is in $\mathbf{b^{\prime}}$ and takes the recovery action same as the expert. Since the learner now behaves like $\pi^{\mathrm{E}}$ on all states in $\mathbf{E}$ and can successfully recovers in $\mathbf{B^{\prime}}$, its total return is the same as the expert, which concludes the proof.
Lemma 23 (Coverage of expert with $N_{\text{off}}$ trajectories).
@@
Recall that we denote by $\mathbf{E}_{\text{annotated}}$ the set of states in $\mathbf{E}$ visited and annotated by offline expert demonstrations. Fix any $s\in\mathbf{E}$ and concatenate all expert trajectories into a single sequence $\{s_{t}\}_{t=1}^{T}$ of length $T:=HN_{\mathrm{off}}$, ordered from the first state of the first trajectory to the last state of the last trajectory. Let $\mathcal{F}_{t}$ be the sigma-algebra generated by all states $s_{1},\dots,s_{t}$, and define the set of initial indices
@@
so by induction,
$\Pr(A_{T})\leq\Big(1-\frac{1-\beta}{N_{0}}\Big)^{T}\leq\exp\!\Big(-\frac{(1-\beta)T}{N_{0}}\Big)=\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big).$
Therefore,
$\Pr\big(s\notin\mathbf{E}_{\text{annotated}}\big)=\Pr(A_{T})\leq\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big).$
@@
$\Pr(\exists s\in\mathbf{E},s\notin\mathbf{E}_{\text{annotated}})\leq N_{0}\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big),$
@@
Appendix E Additional Guarantees for DAgger Variant Without Recoverability Assumption
In this section, we revisit and conduct a refined analysis of another variant of DAgger with trajectory-wise annotations. We show that without the recoverability assumption, an interactive IL algorithm has sample complexity no worse than that of behavior cloning. This result complements prior work [foster2024behavior] that analyzes a different version of DAgger, which they proved to have a worse sample complexity guarantee than behavior cloning.
@@ E.1 Additional Notations and Useful Distance Measures
In line with [foster2024behavior], we consider another oracle that models interacting with the demonstration expert: the trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ that takes into a state-sequence $s_{1:H}$ and returns $a^{*}_{1:H}\sim\pi^{\mathrm{E}}(\cdot\parallel s_{1:H})$. Different from other sections, in the following we adopt Markovian policy, which is a collection of $H$ mappings from states to probability distributions over actions $\pi=\{\pi_{h}:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})\}_{h=1}^{H}$.
We here provide a formal definition of first-step policy mixing—used in the definition of $\hat{\pi}$ (see Algorithm 1, Algorithm 2), which is a common technique (e.g., [sun2017deeply, syed2008apprenticeship, osband2016deep]), defined as follows:
@@
$\Pi_{\mathcal{B}}:=\left\{\pi_{u}:u\in\Delta(\mathcal{B})\right\}$, where policy $\pi_{u}$ is executed in an an episode of an MDP $\mathcal{M}$ by: draw $\pi\sim u$ at the beginning of the episode, and execute policy $\pi$ throughout the episode.
Importantly, $\pi_{u}$ is not a stationary policy (note its difference with $\bar{\pi}_{u}$ in Definition 1); as a result, $a_{1:H}$ are dependent conditioned on $s_{1:H}$, while $a_{1:H}$ are only conditionally independent given $s_{1:H}$ and the random policy $\pi$ drawn.
Additionally, we use $\pi(\cdot\parallel s_{1:H})$ to denote the causally-conditioned probability of action sequence $a_{1:H}$ induced by $\pi$, given state sequence $s_{1:H}$ [ziebart2010modeling].  [^2] The use of $\parallel$ highlights its distinction from standard conditioning on $s_{1:H}$. To elaborate:
  • For Markovian policy $\pi$, $\pi(a_{1:H}\parallel s_{1:H}):=\prod_{h=1}^{H}\pi_{h}(a_{h}|s_{h})$.
It is well-known that the trajectory distribution induced by Markovian policies and their first-step mixings $\pi$ can be factorized to the product of $\pi(a_{1:H}\parallel s_{1:H})$ and the causally-conditioned probability of the state sequence given the action sequence (Definition 40 and Lemma 41). When it is clear from context, we use shorthand $\pi(s_{1:H})$ for $\pi(\cdot\parallel s_{1:H})$.
In the following, we present another 2 useful distance measures for pair of policies.
Definition 25 (Trajectory-wise $L_{\infty}$-semi-metric [foster2024behavior]).
@@
$\rho(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(s_{1:H})}\left[\mathbb{I}\left\{\exists h:a_{h}\neq a^{\prime}_{h}\right\}\right].$
$\rho(\pi\parallel\pi^{\prime})$ is the probability of any action taken by $\pi^{\prime}$ deviating from actions in trajectories induced by $\pi$, which is symmetric [foster2024behavior]. A bound on $\rho(\pi\parallel\pi^{\mathrm{E}})$ leads to straightforward performance difference guarantee: $J(\pi^{\mathrm{E}})-J(\pi)\leq R\cdot\rho(\pi\parallel\pi^{\mathrm{E}})$ [foster2024behavior] (Lemma 37).
@@
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their decoupled Hellinger distance as $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\prime}(s_{1:H}))\right]$.
Similarly, $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(s_{1:H}),\pi^{\prime}(s_{1:H}))\right]$ denotes the expected Hellinger distance between the distribution of actions $\pi(s_{1:H})$ and $\pi^{\prime}(s_{1:H})$ on state sequence $s_{1:H}$ visited by $\pi$. This allows decoupled analysis for state and action sequences, which is useful for the proof of Theorem 27.
E.2 Interactive IL Matches Offline IL on Trajectory-wise Annotation
Next, we consider the trajectory-wise sampling model. We present Tragger, another DAgger variant, namely Algorithm 3 and provide its sample complexity bounds.

### v3

@@ D.2 Lower Bound for Stagger
Throughout the proof, denote by $\mathbf{E}_{\mathrm{annotated}}$ the final set of states in $\mathbf{E}$ on which Stagger have requested expert annotations.
@@
By Lemma 25, if Stagger collects at most $\frac{HN_{0}}{12}$ interactive state-wise annotations as above, then with probability at least $\frac{1}{2}$, $|\mathbf{E}_{\mathrm{annotated}}|$, the number of distinct states in $\mathbf{E}$ annotated is fewer than $N_{0}/3$. Consider a random rollout of $\hat{\pi}$, and define the following events:
$F_{1}$ $:=\left\{s_{1}\notin\mathbf{E}_{\mathrm{annotated}}\right\},$
@@
$\geq$ $\mathbb{P}_{\hat{\pi}}(F_{1},F_{2})\cdot\mathbb{E}_{\hat{\pi}}\left[H-\sum_{h=1}^{H}r_{h}\mid F_{1},F_{2}\right]$
$\geq$ $\frac{2}{3}\cdot\frac{H-8}{H}\cdot\frac{10H-1}{10H}\cdot(H-1)\geq\frac{H}{2},$
where we use our setting of $\beta=\frac{8}{H-8}$ and apply $H\geq 50$. ∎
@@
$\Pr\left(|\mathbf{E}_{\mathrm{annotated}}|\geq\frac{N_{0}}{3}\right)\leq\frac{1}{2}.$
@@
In this proof, we say that state $s$ is annotated at iteration $i$, if it has been annotated by the expert before iteration $i$ (excluding iteration $i$). Since each iteration of Stagger samples one state uniformly from the current policy’s rollout for annotation, we denote the indicator of expert annotating an unannotated state from $\mathbf{E}$ at iteration $i$ as $X_{i}\in\{0,1\}$. With this notation, we have the total number of annotated states by the end of iteration $N_{\mathrm{int}}$ as
$|\mathbf{E}_{\mathrm{annotated}}|=\sum_{i=1}^{N_{\mathrm{int}}}X_{i}.$
Let $\mathcal{F}_{j}$ be the sigma-algebra generated by all information seen by Stagger up to iteration $j$. We now upper bound the expected value of $X_{i}$ conditioned on $\mathcal{F}_{i-1}$ for each $i$.
@@
$\mathbb{E}[|\mathbf{E}_{\mathrm{annotated}}|]=\sum_{i=1}^{N_{\mathrm{int}}}\mathbb{E}[X_{i}]\leq\frac{2N_{\mathrm{int}}}{H}.$
@@
$\Pr(|\mathbf{E}_{\mathrm{annotated}}|\geq N_{0}/3)\leq\frac{\mathbb{E}[X]}{N_{0}/3}\leq\frac{6N_{\mathrm{int}}}{HN_{0}}\leq\frac{1}{2},$
@@ D.3 Upper Bound for Warm-Stagger
We divide the proof into four steps. Throughout, we denote by $\mathbf{E}_{\mathrm{annotated}}$ and $\mathbf{E}_{\mathrm{annotated}}^{\prime}$ the subsets of $\mathbf{E}$ and $\mathbf{E^{\prime}}$, respectively, that are annotated by the $N_{\mathrm{off}}$ offline expert demonstration trajectories.
@@
$F_{3}:=\{\mathbf{E}_{\mathrm{annotated}}=\mathbf{E}\}.$
@@
, we define the event under which only a small fraction of $\mathbf{E^{\prime}}$ is covered by the expert. With our setting of MDP parameters, the number of offline trajectories satisfies
@@
where in the first inequality we apply $H\geq\frac{5}{4}\log(10N_{0})$; in the second inequality, we apply $\frac{H-16}{H-8}<\frac{4}{5}$ when $H\geq 50$; and we apply $N_{0}<\frac{N_{1}}{160}$ in the third inequality by the MDP setting. Thus, by Lemma 21,
$\mathbb{E}\!\left[|\mathbf{E}_{\mathrm{annotated}}^{\prime}|\right]\leq\frac{N_{1}}{20}.$
@@
$\Pr\Big(|\mathbf{E}_{\mathrm{annotated}}^{\prime}|>\frac{N_{1}}{4}-2\Big)\leq\Pr\Big(|\mathbf{E}_{\mathrm{annotated}}^{\prime}|>\frac{N_{1}}{4}-\frac{2N_{1}}{500}\Big)\leq\frac{25}{121}<0.21.$
@@
$F_{4}:=\left\{|\mathbf{E}_{\mathrm{annotated}}^{\prime}|\leq\frac{N_{1}}{4}-2\right\}.$
@@
, we show that under events $F_{3}$ and $F_{4}$, the interactive phase of Warm-Stagger gets $\mathbf{b^{\prime}}$ annotated with probability greater than 0.4 in each of the first three rollouts, such that $\mathbf{b^{\prime}}$ is annotated within three iterations with good probability.
First, we observe that by union bound,
$\Pr(F_{3},F_{4})\geq\Pr(F_{3})+\Pr(F_{4})-1\geq 0.69.$
We henceforth condition on $F_{3}\cap F_{4}$ happening. By the definition of $F_{3}$, the policies $\pi^{n}$ produced by Warm-Stagger acts optimally on all states in $\mathbf{E}$.
Denote by $\mathbf{E}_{\mathrm{annotated}}^{\prime,n}$ the set of annotated states in $\mathbf{E}$ after Warm-Stagger’s iteration $n$; with this notation, $\mathbf{E}_{\mathrm{annotated}}^{\prime,0}=\mathbf{E}_{\mathrm{annotated}}^{\prime}$, and $|\mathbf{E}_{\mathrm{annotated}}^{\prime,n+1}|\leq|\mathbf{E}_{\mathrm{annotated}}^{\prime,n}|+1$ for all $n$. Thus, by the definition of $F_{4}$, for every $n=0,1,2$, $\mathbf{E}_{\mathrm{annotated}}^{\prime,n+1}\leq\frac{N_{1}}{4}-2+2=\frac{N_{1}}{4}$. In other words, at least $\frac{3}{4}$ of the states in $\mathbf{E^{\prime}}$ are unannotated for each of the first three iterations.
During the rollout of a $\pi^{n}$, let $\tau$ be the first step such that $s_{\tau}\in\mathbf{E^{\prime}}$. By Lemma 22, for every $n\in\left\{1,2,3\right\}$,
$\Pr_{\mathcal{M},\pi^{n}}(\tau\leq H/5)\geq 0.79.$
Since by the definition of $\mathcal{M}$, $s_{\tau}$ is drawn uniformly from $\mathbf{E^{\prime}}$ conditioned on $\tau\leq H/5$,
$\Pr_{\mathcal{M},\pi^{n}}(s_{\tau}\notin\mathbf{E}_{\mathrm{annotated}}^{\prime,n}\mid\tau\leq H/5)\geq 0.75.$
Therefore,
$\Pr_{\mathcal{M},\pi^{n}}(s_{\tau}\notin\mathbf{E}_{\mathrm{annotated}}^{\prime,n},\tau\leq H/5)\geq 0.79\times 0.75\geq 0.59$ (11)
@@
$\Pr_{\mathcal{M},\pi^{n}}\big(\forall h\in[H]:s_{h}\in(\mathbf{E}^{\prime}\setminus\mathbf{E}_{\mathrm{annotated}}^{\prime,n})\cup\mathbf{B^{\prime}}\implies\ a_{h}\neq\pi^{E}(s_{h})\big)\;\geq\;0.9.$ (12)
Thus, when the events in Eqs. (11) and  (12) happen simultaneously (which happens with probability $\geq 0.59+0.9-1\geq 0.49$ by union bound), the trajectory rolled out by $\pi^{n}$ transitions to $\mathbf{b^{\prime}}$ at step $\tau+1$ and stays there till the end of episode, accumulating no less than $0.8H$ $\mathbf{b^{\prime}}$ states. Since Warm-Stagger samples one state uniformly from each rollout for annotation, this gives 0.8 probability of $\mathbf{b^{\prime}}$ being annotated with probability $\geq 0.49$.
Denote by $\mathcal{F}_{n}$ the sigma-algebra generated by observations up to iteration $n$ of Warm-Stagger. Our reasoning above implies that for all history $\mathcal{F}_{n-1}$ such that $F_{3}\cap F_{4}$ happens, for all $n\in\left\{1,2,3\right\}$,
$\Pr(\mathbf{b^{\prime}}\text{ is annotated at iteration $n$}\mid\mathcal{F}_{n-1})\geq 0.49\times 0.8\geq 0.39.$
@@
$F_{5}:=\{\mathbf{b^{\prime}}\text{ is annotated within 3 iterations}\}.$
Hence
$\Pr(F_{5}\mid F_{3},F_{4})\;\geq\;1-(1-0.39)^{3}=1-0.61^{3}>0.77.$
@@
, on $F_{3}\cap F_{4}\cap F_{5}$, all states in $\mathbf{E}$ and the reset state $\mathbf{b^{\prime}}$ are annotated, under reward function $R_{1}$, the learner receives reward 1 at any step if it is in $\mathbf{E}$ or $\mathbf{E^{\prime}}$ or it is in $\mathbf{b^{\prime}}$ and takes the recovery action same as the expert. Since learned policy now behaves identical to $\pi^{\mathrm{E}}$ on all states in $\mathbf{E}$ and can successfully recover in $\mathbf{B^{\prime}}$, its total return is the same as the expert, which concludes the proof. ∎
Lemma 27 (Coverage of states in $\mathbf{E}$ with $N_{\mathrm{off}}$ trajectories).
@@
Recall that we denote by $\mathbf{E}_{\mathrm{annotated}}$ the set of states in $\mathbf{E}$ visited and annotated by offline expert demonstrations. Fix any $s\in\mathbf{E}$ and concatenate all expert trajectories into a single sequence $\{s_{t}\}_{t=1}^{T}$ of length $T:=HN_{\mathrm{off}}$, ordered from the first state of the first trajectory to the last state of the last trajectory. Let $\mathcal{F}_{t}$ be the sigma-algebra generated by all states $s_{1},\dots,s_{t}$, and define the set of initial indices
@@
Thus,
$\Pr\big(s\notin\mathbf{E}_{\mathrm{annotated}}\big)=\Pr(A_{T})\leq\Big(1-\frac{1-\beta}{N_{0}}\Big)^{T}\leq\exp\!\Big(-\frac{(1-\beta)T}{N_{0}}\Big)=\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big).$
@@
$\Pr(\exists s\in\mathbf{E},s\notin\mathbf{E}_{\mathrm{annotated}})\leq N_{0}\exp\!\Big(-\frac{(1-\beta)HN_{\mathrm{off}}}{N_{0}}\Big),$
@@
$N_{\mathrm{off}}\;\geq\;\frac{N_{0}}{(1-\beta)H}\log\!\Big(\frac{N_{0}}{\delta}\Big).\qed$
Appendix E Additional Guarantees for a Trajectory-wise DAgger Variant without Recoverability Assumption
In this section, we revisit and conduct a refined analysis of another variant of DAgger with trajectory-wise annotations. We show that without the recoverability assumption, an interactive IL algorithm has sample complexity no worse than that of behavior cloning. This result complements [17] that analyzes a different version of trajectory-wise DAgger, which they proved to have a worse sample complexity guarantee than behavior cloning.
@@ E.1 Additional Notations and Useful Distance Measures
Recall that we have defined first-step mixing of policies in Section C.1. We instantiate it to define a mixture of policies in $\mathcal{B}$, which induces a useful policy class below:
@@
Define $\Pi_{\mathcal{B}}:=\left\{\pi_{u}:u\in\Delta(\mathcal{B})\right\}$, where policy $\pi_{u}$ is executed in an an episode of an MDP $\mathcal{M}$ by: draw $\pi\sim u$ at the beginning of the episode, and execute policy $\pi$ throughout the episode.
In the following, we present another two useful distance measures for a pair of policies.
Definition 29 (Trajectory-wise $L_{\infty}$-semi-metric [17]).
@@
$\rho(\pi\parallel\pi^{\prime}):=\mathbb{E}^{\pi}\mathbb{E}_{a^{\prime}_{1:H}\sim\pi^{\prime}(\cdot\|s_{1:H})}\left[\mathbb{I}\left\{\exists h:a_{h}\neq a^{\prime}_{h}\right\}\right].$
$\rho(\pi\parallel\pi^{\prime})$ is the probability of any action taken by $\pi^{\prime}$ deviating from actions in trajectories induced by $\pi$, which is symmetric [17]. A bound on $\rho(\pi\parallel\pi^{\mathrm{E}})$ leads to straightforward performance difference guarantee: $J(\pi^{\mathrm{E}})-J(\pi)\leq R\cdot\rho(\pi\parallel\pi^{\mathrm{E}})$ [17] (Lemma 40).
Recall that we have defined causually conditional probabilities in Section C.1. Built upon it, we introduce the following definition:
@@
For a pair of Markovian policies $\pi$ and $\pi^{\prime}$, define their decoupled Hellinger distance as $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\|s_{1:H}),\pi^{\prime}(\cdot\|s_{1:H}))\right]$.
Similarly, $\mathbb{E}^{\pi}\left[D^{2}_{\text{H}}(\pi(\cdot\|s_{1:H}),\pi^{\prime}(\cdot\|s_{1:H}))\right]$ denotes the expected Hellinger distance between the causal distribution of actions $\pi(\cdot\|s_{1:H})$ and $\pi^{\prime}(\cdot\|s_{1:H})$ on state sequence $s_{1:H}$ visited by $\pi$. This allows decoupled analysis for state and action sequences, which is useful for our analysis below.
E.2 Interactive IL Matches Offline IL on Trajectory-wise Annotation Sample Complexity
We present Tragger (Algorithm 3), another variant of DAgger that operates in the trajectory-wise sampling model, and provide its sample complexity bounds.